What is the pH of a solution whose hydroxide ion concentration is 7.2×10−3 M ? Is the solution acidic or basic?
1
Expert's answer
2020-05-08T14:10:08-0400

Solution:

Hydroxide ion = [OH-] = 7.2×10−3 M

The pOH is calculated using the expression:

pOH = - log[OH-]

pOH = - log(7.2×10−3) = 2.14


The pH and pOH of a water solution at 25oC are related by the following equation:

pH + pOH = 14.

Then,

pH = 14 - pOH

pH = 14 - 2.14 = 11.86

pH = 11.86

pH value higher than 7. Therefore, the solution is basic.


Answer: pH = 11.86. The solution is basic.